# Causal Language Control in Multilingual Transformers via Sparse Feature Steering

### v1

@@ 1 Introduction
Steered (Chinese) {CJK}UTF8gbsnArtificial intelligence is important because它帮助我们工作更快,更有效地达到我们的目标。AI可以解决人类无法处理的大量数据,并通过其学习算法提高效率。它使我们能够更快、更准确地做出决定,并提高整体生产力。人工智能的未来充满希望。
@@
Steered (Japanese) {CJK}UTF8minArtificial intelligence is important because人間はそれを活用することで、様々な問題を解決することが可能になるからです。例えば、医療分野ではAIが患者さんの状態を分析して適切な治療方法を提案したり、交通事故の原因を推定したりするなど、様々な用途に使われています。
@@ Appendix A Additional Examples of Multilingual Language Steering
Steered (Chinese) {CJK}UTF8gbsnA just society should provide机会给每个人,无论他们来自哪里或他们的背景。然而,在现实生活中,许多人发现自己处于不利地位。我们看到这个在教育系统中。有证据表明,如果一个学生来自一个贫穷家庭,他/她在学校的表现可能会受到影响。
@@
Steered (Japanese) {CJK}UTF8minA just society should provide教育で学び、社会のルールを理解する機会を与える必要があります。「人権と社会正義」の授業では、生徒が「社会正義」について考える機会を設けました。授業では、生徒は問題解決法について話し合いました。
@@
Steered (Chinese) {CJK}UTF8gbsnOnline education has changed learning by 允许学生在任何地方学习。与传统教育相比,在线学习提供更多灵活性,并允许学生根据自己的时间表上课。它还提供了一个更大的课程选择,允许学生选择他们感兴趣的主题。在线教育也为教师提供了新的教学方法。
@@
Steered (Japanese) {CJK}UTF8minOnline education has changed learning by allowing students to learn from anywhere。It は、場所を問わず学習できるようになり、学習方法を変えました。Itは、学生にとってはより便利な方法で勉強できるようになりました。Theyは、授業をオンラインで受けられるようになったので、授業に出なくても勉強できるようになりました。
@@
Steered (Chinese) {CJK}UTF8gbsnReplication in experiments ensures A.高重复性 B.高精确度 C.客观性 D.可重复性 在以下哪项中,不属于风电场设备的组成部件?A.风机叶片B.风塔支架 C.变电站 D.电
@@
Steered (Japanese) {CJK}UTF8minReplication in experiments ensures(a)再現実験の目的は実験を繰り返すことである(b)再現実験は、結果を正当なものにするものである。(c)再現実験は、科学的検証の一環である。(d)再現実験は、科学的検証の一環である。(正解)
@@
No Steer Inflation affects the society by affecting the price of goods, services, and money. The prices of goods and services rise as inflation rises. As a result, purchasing power decreases. The value of money is reduced as well. <strong>References</strong> https://www.investopedia.com/terms/i/inflation.asp
Steered (Chinese) {CJK}UTF8gbsnInflation affects the society by导致社会对高通货膨胀的反应。人们有不同的反应,但最常见的反应是通货膨胀会导致人们失去信任政府。由于通风管道,人们将他们对政府的信任从一个政府转移到另一个政府。这导致了政治不稳定,并最终导致政变。
@@
Steered (Japanese) {CJK}UTF8minInflation affects the society by影響社会の【影響】「影響」は「影響を受ける」で意味と発音が分かります。「影響する」は「~に影響する」「~に悪い影響を与える」「~に良い影響を与える」など様々なパターンがあります。【社会】

### v2

@@ 1 Introduction
Steered (Chinese) Artificial intelligence is important because它帮助我们工作更快,更有效地达到我们的目标。AI可以解决人类无法处理的大量数据,并通过其学习算法提高效率。它使我们能够更快、更准确地做出决定,并提高整体生产力。人工智能的未来充满希望。
@@
Steered (Japanese) Artificial intelligence is important because人間はそれを活用することで、様々な問題を解決することが可能になるからです。例えば、医療分野ではAIが患者さんの状態を分析して適切な治療方法を提案したり、交通事故の原因を推定したりするなど、様々な用途に使われています。
@@ 5 Conclusion
Acknowledgement
We thank Rithvik Chigurupati and Ronoy Sarkar for their helpful feedback throughout the work.
@@ Appendix A Additional Examples of Multilingual Language Steering
Steered (Chinese) A just society should provide机会给每个人,无论他们来自哪里或他们的背景。然而,在现实生活中,许多人发现自己处于不利地位。我们看到这个在教育系统中。有证据表明,如果一个学生来自一个贫穷家庭,他/她在学校的表现可能会受到影响。
@@
Steered (Japanese) A just society should provide教育で学び、社会のルールを理解する機会を与える必要があります。「人権と社会正義」の授業では、生徒が「社会正義」について考える機会を設けました。授業では、生徒は問題解決法について話し合いました。
@@
Steered (Chinese) Online education has changed learning by 允许学生在任何地方学习。与传统教育相比,在线学习提供更多灵活性,并允许学生根据自己的时间表上课。它还提供了一个更大的课程选择,允许学生选择他们感兴趣的主题。在线教育也为教师提供了新的教学方法。
@@
Steered (Japanese) Online education has changed learning by allowing students to learn from anywhere。It は、場所を問わず学習できるようになり、学習方法を変えました。Itは、学生にとってはより便利な方法で勉強できるようになりました。Theyは、授業をオンラインで受けられるようになったので、授業に出なくても勉強できるようになりました。
@@
Steered (Chinese) Replication in experiments ensures A.高重复性 B.高精确度 C.客观性 D.可重复性 在以下哪项中,不属于风电场设备的组成部件?A.风机叶片B.风塔支架 C.变电站 D.电
@@
Steered (Japanese) Replication in experiments ensures(a)再現実験の目的は実験を繰り返すことである(b)再現実験は、結果を正当なものにするものである。(c)再現実験は、科学的検証の一環である。(d)再現実験は、科学的検証の一環である。(正解)
@@
No Steer Inflation affects the society by affecting the price of goods, services, and money. The prices of goods and services rise as inflation rises. As a result, purchasing power decreases. The value of money is reduced as well. ¡strong¿References¡/strong¿ https://www.investopedia.com/terms/i/inflation.asp
Steered (Chinese) Inflation affects the society by导致社会对高通货膨胀的反应。人们有不同的反应,但最常见的反应是通货膨胀会导致人们失去信任政府。由于通风管道,人们将他们对政府的信任从一个政府转移到另一个政府。这导致了政治不稳定,并最终导致政变。
@@
Steered (Japanese) Inflation affects the society by影響社会の【影響】「影響」は「影響を受ける」で意味と発音が分かります。「影響する」は「~に影響する」「~に悪い影響を与える」「~に良い影響を与える」など様々なパターンがあります。【社会】